OpenStudy (surryyy):

Select the equation of the line parallel to the equation y=-4x-6 and that passes through the point (1, 2). y=4x+6 y= - 4x +6 y = -4x - 4 y= 3x -1

OpenStudy (sohailiftikhar):

here you just have to look at the slop of line which matches the given line because slop of parallel is same..

OpenStudy (sohailiftikhar):

so what you think know ? what's the answer ? :)

OpenStudy (sohailiftikhar):

this is the general equation y=mx+c here the coefficient of x is called slop ..

OpenStudy (sohailiftikhar):

the slop of given line is -4

OpenStudy (surryyy):

would it be b?

OpenStudy (sohailiftikhar):

now as there are two options with slop=-4 so we use the general formula to find equation by using given pionts x1=1 and y1=2 y-y1/x-x1=-4 y-2/x-1=-4 y-2=-4x+4 y=-4x+6 yes it's b

OpenStudy (surryyy):

thnx

OpenStudy (sohailiftikhar):

np